A company is considering an acquisition of The Company “B”. B has a cost of equity...

  1. A company is considering an acquisition of The Company “B”. B has a cost of equity of 10% and 25% of its financing is in the form debt at 6%. After acquisition, it is estimated that the cash flows and the interest payments for the next three years are as follows:

                            Year 1          Year 2          Year 3

       FCF $10 $20 $25

       Interest expense 28                24           20.28

The cash flows are then expected to grow at a constant rate of 5% indefinitely.       

What is the value of the operations to the acquiring company?
Answer #1

Weight of Debt (wd)= 0.25

Weight of Equity (we)= 0.75

Cost of equity (ce)= 0.10

Pre-tax cost of debt (cd)= 0.06

Post tax cost of debt assuming tax is nil is 0.06

Weighted average cost of capital= (we*ce)+(wd*cd*(1-t)) here t=0

(0.75*0.10)+(0.25*0.06)= 0.09=9%

Free cash flow to firm (FCFF)= FCF

FCF= EBITDA-Fixed Capital Investment- change in Working capital+Non Cash expenses

Assumptions

1) T=0

So EAT=EBT

2) D&A=0. So NCC=0

EBIT= EBITDA

Therefore FCFF= EBIT-Interest+non cash charges+Int(1-tax rate)-Fixed Capital Investment-Change in Working capital

FCFF=FCF

Terminal Value after 3rd year is ((25*(1+g))/(r-g))

g=0.05

r=cost of capital=0.09

Terminal value is calculated to 3rd year cash flow

Particulars Year 1 Year 2 Year 3
FCF 10 20 25
FCFF 10 20 550
Discount factors 1.09 1.1881 1.295029
Present value (Respective year FCFF/Discount factor) 9.174312 16.8336 424.7009
Sum of present value 450.7088

Value of operations= $450.7088